This question is typical on some driver’s li- cense exams: A car moving at 47 km/h skids 12 m with locked brakes.
How far will the car skid with locked brakes at 94 km/h? Assume that energy loss is due only to sliding friction.
Answer in units of m.
Answer #1

Since the road is same, So coefficient of friction will also be same, which means de-acceleration during brake will be same in both cases, Using this in 3rd kinematic equation:

V^2 = U^2 + 2*a*S

Since V = final speed = 0 m/sec

and 'a' is negative and constant, So

U^2 = 2*a*S

S = distance travelled during skidding will be proportional to U^2

S2/S1 = (U2/U1)^2

S2 = S1*(U2/U1)^2

S1 = 12 m

U2 = 94 km/hr

U1 = 47 km/hr

Using these values:

S2 = 12*(94/47)^2

S2 = 48 m

Please Upvote.